Labor Force Couples The percentage of couples where both parties are in the labor force is 52.1. Choose 5 couples at random. Find the probability that
a. None of the couples have both persons working
b. More than 3 of the couples have both persons in the labor force
c. Fewer than 2 of the couples have both parties working


Answer by Boreal(15235)

the first is p(0) is (1-p)^5=0.479^5=0.0252
more than 3 means 4 or 5
for p(4)=5C4*0.521^4*0.479^1, binomial formula. This equals 0.1765
for p(5)=0.521^5=0.0384
The answer is the sum or 0.2149
--
p(1) is 5*0.521*0.479^4=0.1371
added to p(0) gives the probability of fewer than 2, and that is 0.1623.